# Towards an Explainable Comparison and Alignment of Feature Embeddings

### v2

@@ Abstract
While several feature embedding models have been developed in the literature, comparisons of these embeddings have largely focused on their numerical performance in classification-related downstream applications. However, an interpretable comparison of different embeddings requires identifying and analyzing mismatches between sample groups clustered within the embedding spaces. In this work, we propose the Spectral Pairwise Embedding Comparison (SPEC) framework to compare embeddings and identify their differences in clustering a reference dataset. Our approach examines the kernel matrices derived from two embeddings and leverages the eigendecomposition of the difference kernel matrix to detect sample clusters that are captured differently by the two embeddings. We present a scalable implementation of this kernel-based approach, with computational complexity that grows linearly with the sample size. Furthermore, we introduce an optimization problem using this framework to align two embeddings, ensuring that clusters identified in one embedding are also captured in the other model. We provide numerical results demonstrating the SPEC’s application to compare and align embeddings on large-scale datasets such as ImageNet and MS-COCO. The project page is available at https://mjalali.github.io/SPEC/.
@@ 1 Introduction
Several mainstream frameworks in computer vision and natural language processing rely on embedding models to map raw image and text inputs into spaces with semantically meaningful features (Radford et al., 2021; Liu et al., 2023; Alayrac et al., 2022; Li et al., 2023). The application of pre-trained embeddings has enabled scalable solutions for many downstream tasks, particularly in scenarios where the available sample size and compute resources are significantly limited. In such cases, the embedded data can be used to train simple models, such as linear or k-nearest neighbors (KNN) classifiers, to achieve satisfactory results. Additionally, features extracted by standard embedding models are widely employed for the automated evaluation of generative models (Heusel et al., 2017; Kynkäänniemi et al., 2023; Stein et al., 2023), providing accurate rankings of generative modeling architectures without requiring time-intensive human assessments.
While recent advancements in the machine learning community have introduced various embedding models that achieve remarkable results on standard image, text, and video domains, comparisons of these embeddings have primarily focused on evaluating their performance in standard downstream tasks, such as classification accuracy on benchmark datasets (e.g. ImageNet). However, such comparisons often lack interpretability and do not reveal how differently the embeddings behave in recognizing various sample types (Boggust et al., 2022). A more fine-grained comparison is necessary to disclose explainable differences between embedding models, particularly in identifying which samples are clustered differently according to the models. Understanding these differences can aid in interpreting and debugging embeddings and can also be leveraged to align multiple embeddings (Simhi & Markovitch, 2023; Dar et al., 2023). Furthermore, interpreting the discrepancies between embeddings can be utilized to select representation models for downstream applications such as generative model evaluation (Stein et al., 2023; Kynkäänniemi et al., 2023).
@@
In this work, we propose a spectral approach called Spectral Pairwise Embedding Comparison (SPEC) for the fine-grained comparison of two embeddings. The SPEC framework detects differences in sample clusters assigned by two embeddings, identifying major data groups that are clustered differently by one embedding compared to the other one. To achieve this, we adopt standard spectral clustering, which leverages the eigendecomposition of the kernel similarity matrix, and propose analyzing the principal eigenvectors of the difference of embeddings’ kernel matrices to interpret the differences in cluster assignments. Our analysis suggests that the SPEC framework can effectively detect cluster differences between two embeddings.
To address the computational challenges of performing eigendecomposition on large-scale datasets, we develop a scalable implementation of SPEC. A direct eigendecomposition of the $n\times n$ difference kernel matrix requires $O(n^{3})$ computations for a dataset with $n$ samples, which is computationally expensive for large datasets. Assuming a bounded feature dimension $d$ for the applied kernel function, we prove that the eigenspace of the difference kernel matrix can be computed using $O(\max\{d^{3},n\})$ operations, resulting in a scalable algorithm under a moderate dimension $d$ value. Furthermore, we extend this scalable computation method to shift-invariant kernel functions, e.g. the Gaussian kernel, by employing the framework of random Fourier features (RFF) (Rahimi & Recht, 2007a), where the size of RFF proxy-feature map can be controlled for a more efficient application of SPEC.
We also explore the application of the SPEC framework to define a distance measure between two embeddings. We define the $\mathrm{SPEC}$-$\mathrm{diff}$ distance as the spectral radius of the kernel difference matrix, which aims to quantify the weight of the most differently captured cluster in one embedding that is not strongly clustered by the other model. We discuss scalable computations of this distance and its gradient with respect to the embedding parameters. Using the power method and the calculated left and right eigenvectors of the differential covariance matrix, we enable gradient-based optimization of the distance measure for aligning embedding models. This gradient-based approach leads to a method we call SPEC-align, aligning embedding models by minimizing their differences in clustering a reference dataset. SPEC-align is particularly useful for aligning cross-modality embeddings, such as CLIP (Radford et al., 2021), with a state-of-the-art single-modality embedding. Such a spectral alignment can improve the performance of cross-modality embeddings in capturing concepts specific to individual modalities.
@@ 2 Related Work
Spectral Clustering, Kernel PCA, and Random Fourier Features. Kernel PCA (Schölkopf et al., 1998) is a widely recognized technique for dimensionality reduction that relies on the eigendecomposition of the kernel matrix. Several studies (Bengio et al., 2003b, a) have explored the relationship between kernel PCA and spectral clustering. Also, the analysis of random Fourier features (Rahimi & Recht, 2007b) for performing scalable kernel PCA has been studied by Chitta et al. (2012); Ghashami et al. (2016); Ullah et al. (2018); Sriperumbudur & Sterge (2022); Gedon et al. (2023). In this paper, we introduce a spectral approach for comparing two embeddings, leveraging the random Fourier features framework to address computational challenges. Unlike Laplacian spectral clustering (Ng et al., 2001) that uses the graph Laplacian, our method uses the kernel matrix similar to Kernel PCA.
@@
Darrin et al. (2024) propose a new metric for comparing embeddings without labeled data and propose the concept of information sufficiency (IS) to quantify the required information to simulate one embedding from another. Our work offers a complementary, explainable method for comparing embeddings by detecting different sample clusters assigned by embeddings and providing a method for aligning them.
A different yet related line of work is the evaluation of generative models. (Bińkowski et al., 2018; Jalali et al., 2023; Ospanov et al., 2024; Jalali et al., 2024; Ospanov & Farnia, 2024) leverage the eigenspectrum of kernel matrices to quantify diversity. The papers (Jiralerspong et al., 2023; Zhang et al., 2024) explore novelty evaluation, analyzing how generated samples differ from those of a reference distribution. In particular, (Zhang et al., 2024, 2025) propose a spectral method for measuring the entropy of the novel modes of a generative model with respect to a reference model, relying on the eigendecomposition of kernel similarity matrices.
Embeddings Alignment. There are many works on embedding alignment for multimodal models (Bellagente et al., 2023; Lu et al., 2024; Han et al., 2024; Wang et al., 2023b; Girdhar et al., 2023; Grave et al., 2019). Salman et al. (2024); Eslami & de Melo (2025) introduced a method, which demonstrates that adversarial perturbations can force text embeddings to align with any image in multimodal models, exposing security vulnerabilities in vision-language learning. Ye et al. (2024) proposed ModalChorus, an interactive system that visualizes and corrects misalignments in multi-modal embeddings, improving interpretability and optimization. Focusing on fine-grained alignment, Yin et al. (2024) introduced a method for explicitly aligning individual word embeddings with corresponding visual features, leveraging cross-modal attention to refine token-image associations. In contrast, our work focuses on aligning embeddings in a kernel setting specifically to match their sample clusters, leading to a different approach to embedding comparison.
@@ 3.2 Kernel Functions and Covariance Matrix
Well-known examples of kernel functions include the cosine-similarity kernel $k_{\text{cosine}}(x,y)=\frac{x^{\top}y}{\|x\|_{2}\|y\|_{2}}$ and the Gaussian (RBF) kernel defined for bandwidth $\sigma$ as:
$k(x,y)=\exp\left(\frac{-\|x-y\|^{2}_{2}}{2\sigma^{2}}\right)$
@@ 4 SPEC: A Spectral Identification of Embeddings’ Mismatches
To compare the two embeddings, note that their corresponding spaces $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ may have different dimensions, and therefore a sample-specific comparison of the embedded vectors $\psi_{1}(x),\,\psi_{2}(x)$ for each individual data point $x$ will not provide a meaningful comparison of the embedding maps. Therefore, a more relevant approach is to compare the embeddings’ outputs over the entire set of data $\{x_{1},\ldots,x_{n}\}$ and investigate which structures are dissimilar between the sets of embedded data following the embeddings. Here, we consider a spectral approach and particularly focus on the difference of kernel matrices between the two embeddings. In the following, we discuss how the eigenspace of the kernel difference matrix can help identify the differently clustered points by the two embeddings.
To do this, consider the kernel matrix of the first embedding $K_{\psi_{1}}=\bigl{[}k_{1}(\psi_{1}(x_{i}),\psi_{1}(x_{j}))\bigr{]}_{(i,j)=(1,1)}^{(n,n)}$ and the kernel matrix of the second embedding $K_{\psi_{2}}=\bigl{[}k_{2}(\psi_{2}(x_{i}),\psi_{2}(x_{j}))\bigr{]}_{(i,j)=(1,1)}^{(n,n)}$.
@@
We propose the framework of Spectral Pairwise Embedding Comparison (SPEC) where the two embeddings $\psi_{1}$ and $\psi_{2}$ are compared using the eigendirections of the difference kernel matrix $\Lambda_{\psi_{1},\psi_{2}}$. As we will show, the principal eigenvectors can be interpreted as the clusters of samples assigned by embedding $\psi_{1}$ that are less strongly grouped by the second embedding $\psi_{2}$. In what follows, we first show a theoretical result supporting the mentioned property of $\Lambda_{\psi_{1},\psi_{2}}$’s eigenvectors. Next, we provide a scalable computation method for computing the eigenspace of $\Lambda_{\psi_{1},\psi_{2}}$ that linearly scales with the sample size $n$.
@@
  • Condition 1: Suppose the sample set $X_{\mathcal{I}}$ characterized by index set $\mathcal{I}$ are separated from the rest of samples by embedding $\psi_{1}$, where the normalized block kernel matrix $\frac{1}{n}K_{\psi_{1}}[\mathcal{I},\mathcal{I}^{c}]$ has a bounded Frobenius norm $\bigl{\|}\frac{1}{n}K_{\psi_{1}}[\mathcal{I},\mathcal{I}^{c}]\bigr{\|}_{F}\leq\epsilon_{1}$.
  • Condition 2: Suppose the sample set $X_{\mathcal{I}}$ characterized by index set $\mathcal{I}$ are weakly grouped by embedding $\psi_{2}$, where the normalized block kernel matrix $\frac{1}{n}K_{\psi_{2}}[\mathcal{I},\mathcal{I}^{c}]$ has a bounded $\ell_{2}$-operator norm (or maximum eigenvalue for this PSD matrix) $\bigl{\|}\frac{1}{n}K_{\psi_{2}}[\mathcal{I},\mathcal{I}]\bigr{\|}_{2}\leq\epsilon_{2}$.
@@
Consider the difference kernel matrix $\Lambda_{\psi_{1},\psi_{2}}$ in (3). Suppose Conditions 1 and 2 hold. Let $\mathbf{v}_{1},\ldots,\mathbf{v}_{n}$ be the unit-norm eigenvectors of $\Lambda_{\psi_{1},\psi_{2}}$ corresponding to eigenvalues $\lambda_{1},\ldots,\lambda_{n}$. For every $i\in\{1,\ldots,n\}$, we define $\lambda^{\mathcal{I}}_{i}$ and $\lambda^{\mathcal{I}^{c}}_{i}$ to be the closest eigenvalue of $\Lambda_{\psi_{1},\psi_{2}}[\mathcal{I},\mathcal{I}]$ and $\Lambda_{\psi_{1},\psi_{2}}[\mathcal{I}^{c},\mathcal{I}^{c}]$ to $\lambda_{i}$. Then, the following holds for $\xi=4(\epsilon_{1}^{2}+\epsilon_{2})$:
@@
The above corollary proves that if an eigenvalue $\lambda$ of the kernel difference matrix $\Lambda_{\psi_{1},\psi_{2}}$ is sufficiently large, such that its gap with the maximum eigenvalue of the block $\Lambda_{\psi_{1},\psi_{2}}[\mathcal{I}^{c},\mathcal{I}^{c}]$ (with the complement of samples in $\mathcal{I}$ clustered by $\psi_{1}$ yet not by $\psi_{2}$) is higher than the threshold $\lambda$, then the $\mathcal{I}^{c}$-entries of the corresponding unit-norm eigenvector $\mathbf{v}$ will be bounded, reflecting the lack of $\mathcal{I}^{c}$ samples in the differentially clustered samples by embedding $\psi_{1}$ and $\psi_{2}$. Based on the above theoretical results, we propose considering the principal eigendirections of the difference kernel matrix, and using their significant-value entries to find the subset of samples clustered by embedding $\psi_{1}$ but not grouped by $\psi_{2}$.
Since the difference kernel matrix is of size $n\times n$, a standard eigendecomposition will cost $O(n^{3})$ computations. Proposition 1 shows that the computation cost will be lower for embeddings with bounded feature maps. In fact, this result shows the computation of the eigenspace can be performed using linearly growing computation cost $O(n)$.
@@
Consider the difference kernel matrix $\Lambda_{\psi_{1},\psi_{2}}$ in (3). This matrix shares the same non-zero eigenvalues with the following matrix:
$\Gamma_{\psi_{1},\psi_{2}}=\begin{bmatrix}C_{\psi_{1}}&C_{\psi_{1},\psi_{2}}\vspace{2mm}\\ -C^{\top}_{\psi_{1},\psi_{2}}&-C_{\psi_{2}}\end{bmatrix}\in\mathbb{R}^{(d_{1}+d_{2})\times(d_{1}+d_{2})}$ (4)
@@
We also note that for every eigenvector $\mathbf{v}\in\mathbb{R}^{d_{1}+d_{2}}$ of the matrix $\Gamma_{\psi_{1},\psi_{2}}$ in (4), which we call the differential covariance matrix, we can find the corresponding vector $\mathbf{u}$ of difference kernel matrix $\Lambda_{\psi_{1},\psi_{2}}$ using the following:
@@
As can be seen, the computation of the matrix $\Gamma_{\psi_{1},\psi_{2}}$ can be performed with $O(n)$ computations, linearly growing in sample size, and the eigendecomposition of $\Gamma_{\psi_{1},\psi_{2}}$ can be handled via $O\bigl{(}(d_{1}+d_{2})^{3}\bigr{)}$, depending on the dimensions of the kernel feature maps $\phi_{1},\phi_{2}$, and finally the eigenvector mapping from $\Gamma_{\psi_{1},\psi_{2}}$ to $\Lambda_{\psi_{1},\psi_{2}}$ will be $O(n)$. Therefore, the entire eigenvector computation of $\Lambda_{\psi_{1},\psi_{2}}$ can be handled using $O(n+(d_{1}+d_{2})^{3})$ computations. Algorithm 1 contains the main steps of computing the SPEC-eigendirections using the above approach. As detailed in this algorithm, the computation of the differential kernel covariance matrix can be run over samples in a cascade, avoiding the need for storing a large dataset.
Applying the standard linear and cosine-similarity kernels, the kernel feature dimension will match that of the embedding, which is usually bounded by 1000 for standard image and text embeddings. In the case of shift-invariant kernels, e.g. the Gaussian (RBF) kernel, whose feature dimension is infinite, we can leverage the random Fourier features (RFFs) (Rahimi & Recht, 2007b) to reduce the dimension of the kernel feature dimension for a proper proxy kernel function characterized by the random Fourier features. According to the RFF framework, given a kernel function $k(x,y)=\kappa(x-y)$ that is normalized i.e. $\kappa(0)=1$, we draw a number $m$ independent Fourier features $\omega_{1},\ldots,\omega_{m}\sim\widehat{\kappa}$ from probability density function $\widehat{\kappa}$ which denotes the Fourier transform of $\kappa$ defined as
@@
Then, the RFF method approximates the shift-invariant kernel $k(x,y)\approx\widehat{k}(x,y)=\widehat{\phi}(x)^{\top}\widehat{\phi}(y)$ where
@@
6:     Update $C_{\psi_{1},\psi_{2}}\leftarrow C_{\psi_{1},\psi_{2}}+\frac{1}{n}\phi_{1}(\psi_{1}(\mathbf{x}_{i}))\phi_{2}(\psi_{2}(\mathbf{x}_{i}))^{\top}$

### v3

@@ Abstract
While several feature embedding models have been developed in the literature, comparisons of these embeddings have largely focused on their numerical performance in classification-related downstream applications. However, an interpretable comparison of different embeddings requires identifying and analyzing mismatches between sample groups clustered within the embedding spaces. In this work, we propose the Spectral Pairwise Embedding Comparison (SPEC) framework to compare embeddings and identify their differences in clustering a reference dataset. Our approach examines the kernel matrices derived from two embeddings and leverages the eigendecomposition of the kernel difference matrix to detect sample clusters that are captured differently by the two embeddings. We present a scalable implementation of this kernel-based approach, with computational complexity that grows linearly with the sample size. Furthermore, we introduce an optimization problem using this framework to align two embeddings, ensuring that clusters identified in one embedding are also captured in the other model. We provide numerical results demonstrating the SPEC’s application to compare and align embeddings on large-scale datasets such as ImageNet and MS-COCO. The project page is available at https://mjalali.github.io/SPEC/.
@@ 1 Introduction
Several mainstream frameworks in computer vision and natural language processing rely on embedding models to map raw image and text inputs into spaces with semantically meaningful features (Radford et al., 2021; Liu et al., 2023; Alayrac et al., 2022; Li et al., 2023; Oquab et al., 2024). The application of pre-trained embeddings has enabled scalable solutions for many downstream tasks, particularly in scenarios where the available sample size and compute resources are significantly limited. In such cases, the embedded data can be used to train simple models, such as linear or k-nearest neighbors (KNN) classifiers, to achieve satisfactory results. Additionally, features extracted by standard embedding models are widely employed for the automated evaluation of generative models (Heusel et al., 2017; Hessel et al., 2021; Stein et al., 2023; Ospanov et al., 2025), providing accurate rankings of generative modeling architectures without requiring time-intensive human assessments.
While recent advancements in the machine learning community have introduced various embedding models that achieve remarkable results on standard image, text, and video domains, comparisons of these embeddings have primarily focused on evaluating their performance in standard downstream tasks, such as classification accuracy on benchmark datasets (e.g. ImageNet). However, such comparisons often lack interpretability and do not reveal how differently the embeddings behave in recognizing various sample types. A more fine-grained comparison is necessary to disclose explainable differences between embedding models, particularly in identifying which samples are clustered differently according to the models. Understanding these differences can aid in interpreting and debugging embeddings and can also be leveraged to align multiple embeddings. Furthermore, interpreting the discrepancies of embeddings can be used to select representation models for downstream tasks.
@@
In this work, we propose a spectral approach called Spectral Pairwise Embedding Comparison (SPEC) for the fine-grained comparison of two embeddings. The SPEC framework detects differences in sample clusters assigned by two embeddings, identifying major data groups that are clustered differently by one embedding compared to the other. To achieve this, we extend the well-established kernel principal component analysis (Kernel-PCA) approach (Schölkopf et al., 1998), that leverages the eigendecomposition of the kernel similarity matrix, and propose analyzing the principal eigenvectors of the difference of embeddings’ kernel matrices to interpret the differences in cluster assignments. Our analysis suggests that the SPEC framework could effectively detect cluster differences between two embeddings, particularly when the clusters can be separated using the principal eigendirections of the kernel matrices.
To address the computational challenges of performing eigendecomposition on large-scale datasets, we develop a scalable implementation of SPEC. A direct eigendecomposition of the $n\times n$ kernel difference matrix requires $O(n^{3})$ computations for a dataset with $n$ samples, which is computationally expensive for large datasets. Assuming a bounded feature dimension $d$ for the applied kernel function, we prove that the eigenspace of the kernel difference matrix can be computed using $O\bigl{(}d^{2}n+d^{3}\bigr{)}$ operations, resulting in a scalable algorithm under a moderate dimension $d$ value. Furthermore, we extend this scalable computation method to shift-invariant kernel functions, e.g. the Gaussian (RBF) kernel, by employing the framework of random Fourier features (RFF) (Rahimi & Recht, 2007), where the size of RFF proxy-feature map can be controlled for a more efficient application of SPEC.
We also explore the application of the SPEC framework to define a distance measure between two embeddings. We define the $\mathrm{SPEC}$-$\mathrm{diff}$ distance as the spectral radius of the kernel difference matrix, which aims to quantify the weight of the most differently captured cluster in one embedding that is not strongly clustered by the other model. We discuss scalable computations of this distance and its gradient with respect to the embedding parameters. Using the power method and the calculated left and right eigenvectors of the differential covariance matrix, we enable gradient-based optimization of the distance measure for aligning embedding models. This gradient-based approach leads to a method we call SPEC-align, aligning embedding models by minimizing their differences in clustering a reference dataset. SPEC-align is particularly useful for aligning cross-modality embeddings, such as CLIP (Radford et al., 2021), with a state-of-the-art single-modality embedding. The alignment can improve the performance of cross-modality embeddings in capturing concepts specific to individual modalities.
@@ 2 Related Work
Spectral Clustering, Kernel PCA, and Random Fourier Features. Kernel PCA (Schölkopf et al., 1998) is a widely established method for dimensionality reduction that relies on the eigendecomposition of the kernel matrix. Several studies (Bengio et al., 2003b, a) have explored the relationship between kernel PCA and spectral clustering. Also, the analysis of random Fourier features (Rahimi & Recht, 2007) for performing scalable kernel PCA has been studied by Chitta et al. (2012); Ghashami et al. (2016); Ullah et al. (2018); Sriperumbudur & Sterge (2022); Gedon et al. (2023). In our work, we introduce a spectral approach for comparing two embeddings, leveraging the random Fourier features framework to address computational challenges.
@@
Darrin et al. (2024) propose a new metric for comparing embeddings without labeled data and propose the concept of information sufficiency (IS) to quantify the required information to simulate one embedding from another. Our work offers a complementary, explainable method for comparing embeddings by detecting different sample clusters assigned by embeddings and providing a method for aligning them. Also, Vargas et al. (2024) discuss a spectral approach to the comparison of embeddings, where they consider the PCA-principal components of the collective embedded vectors of two representations. On the other hand, our work considers the kernel difference matrix, which better suits the comparison of embeddings in different dimensions and leads to a distance metric for their alignment.
A different yet related line of work is the evaluation of generative models. (Bińkowski et al., 2018; Jalali et al., 2023; Ospanov et al., 2024; Jalali et al., 2024; Ospanov & Farnia, 2025; Wang et al., 2023b; Hu et al., 2025a; Rezaei et al., 2025; Hu et al., 2025b, c) leverage the eigenspectrum of kernel matrices to quantify diversity. The papers (Jiralerspong et al., 2023; Zhang et al., 2024) explore novelty evaluation, analyzing how generated samples differ from those of a reference model. In particular, (Zhang et al., 2024, 2025) propose a spectral method for measuring the entropy of the novel modes of a generative model with respect to a reference model.
Embeddings Alignment. There are many works on embedding alignment for multimodal models (Bellagente et al., 2023; Lu et al., 2024; Han et al., 2024; Wang et al., 2023c; Girdhar et al., 2023; Grave et al., 2019). Salman et al. (2024); Eslami & de Melo (2025) introduced a method, which demonstrates that adversarial perturbations can force text embeddings to align with any image in multimodal models, exposing security vulnerabilities in vision-language learning. Ye et al. (2024) proposed ModalChorus, an interactive system that visualizes and corrects misalignments in multi-modal embeddings, improving interpretability and optimization. Focusing on fine-grained alignment, Yin et al. (2024) introduced a method for explicitly aligning individual word embeddings with corresponding visual features, leveraging cross-modal attention to refine token-image associations. In contrast, our work focuses on aligning embeddings in a kernel setting to match their sample clusters.
@@ 3.2 Kernel Functions and Covariance Matrix
Well-known examples of kernel functions include the cosine-similarity kernel $k_{\text{cosine}}(x,y)=\frac{x^{\top}y}{\|x\|_{2}\|y\|_{2}}$ and the Gaussian (RBF) kernel defined for bandwidth $\sigma$ as
$k_{\text{Gaussian}(\sigma^{2})}(x,y)=\exp\Bigl{(}\frac{-\|x-y\|^{2}_{2}}{2\sigma^{2}}\Bigr{)}$
@@ 4 SPEC: A Spectral Identification of Embeddings’ Mismatches
To compare the two embeddings, note that their corresponding spaces $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ may have different dimensions, and therefore a sample-specific comparison of the embedded vectors $\psi_{1}(x),\,\psi_{2}(x)$ for each individual data point $x$ will not provide a meaningful comparison of the embedding maps. Therefore, a more relevant approach is to compare the embeddings’ outputs over the entire set of data $\{x_{1},\ldots,x_{n}\}$ and investigate which structures are dissimilar between the sets of embedded data following the embeddings. Here, we consider a spectral approach and particularly focus on the difference of kernel matrices between the two embeddings. In the following, we discuss how the eigenspace of the kernel difference matrix can help identify the differently clustered points by the two embeddings. In the following, we denote the kernel matrix of the first embedding with $K_{\psi_{1}}=\bigl{[}k_{1}(\psi_{1}(x_{i}),\psi_{1}(x_{j}))\bigr{]}_{(i,j)=(1,1)}^{(n,n)}$ and the kernel matrix of the second embedding with $K_{\psi_{2}}=\bigl{[}k_{2}(\psi_{2}(x_{i}),\psi_{2}(x_{j}))\bigr{]}_{(i,j)=(1,1)}^{(n,n)}$.
@@
We propose the framework of Spectral Pairwise Embedding Comparison (SPEC) where the two embeddings $\psi_{1}$ and $\psi_{2}$ are compared using the eigendirections of the kernel difference matrix $\Lambda_{\psi_{1},\psi_{2}}$. As we will show, the principal eigenvectors can be interpreted as the clusters of samples assigned by embedding $\psi_{1}$ that are less strongly grouped by the second embedding $\psi_{2}$. In what follows, we first show a theoretical result supporting the mentioned property of $\Lambda_{\psi_{1},\psi_{2}}$’s eigenvectors. Next, we provide a scalable computation method for computing the eigenspace of $\Lambda_{\psi_{1},\psi_{2}}$ that linearly scales with the sample size $n$.
@@
  • Condition 1: Suppose the sample set $X_{\mathcal{I}}$ characterized by index set $\mathcal{I}$ are separated from the rest of samples by embedding $\psi_{1}$, where the normalized block kernel matrix $\frac{1}{n}K_{\psi_{1}}[\mathcal{I},\mathcal{I}^{c}]$ has an $\epsilon_{1}$-bounded Frobenius norm:
    $\Bigl{\|}\frac{1}{n}K_{\psi_{1}}\bigl{[}\mathcal{I},\mathcal{I}^{c}\bigr{]}\Bigr{\|}_{F}\leq\epsilon_{1}$
  • Condition 2: Suppose the sample set $X_{\mathcal{I}}$ characterized by index set $\mathcal{I}$ are weakly grouped by embedding $\psi_{2}$, where the normalized block kernel matrix $\frac{1}{n}K_{\psi_{2}}[\mathcal{I},\mathcal{I}]$ has an $\epsilon_{2}$-bounded $\ell_{2}$-operator norm (which is the maximum eigenvalue for this PSD matrix)
    $\Bigl{\|}\frac{1}{n}K_{\psi_{2}}\bigl{[}\mathcal{I},\mathcal{I}\bigr{]}\Bigr{\|}_{2}\leq\epsilon_{2}$
@@
Consider the kernel difference matrix $\Lambda_{\psi_{1},\psi_{2}}$ in (3). Suppose Conditions 1 and 2 hold. Let $\mathbf{v}_{1},\ldots,\mathbf{v}_{n}$ be the unit-norm eigenvectors of $\Lambda_{\psi_{1},\psi_{2}}$ corresponding to eigenvalues $\lambda_{1},\ldots,\lambda_{n}$. For every $i\in\{1,\ldots,n\}$, we define $\lambda^{\mathcal{I}}_{i}$ and $\lambda^{\mathcal{I}^{c}}_{i}$ to be the closest eigenvalue of $\Lambda_{\psi_{1},\psi_{2}}[\mathcal{I},\mathcal{I}]$ and $\Lambda_{\psi_{1},\psi_{2}}[\mathcal{I}^{c},\mathcal{I}^{c}]$ to $\lambda_{i}$. Then, the following holds for $\xi=4(\epsilon_{1}^{2}+\epsilon_{2})$:
@@
The above corollary proves that if an eigenvalue $\lambda$ of the kernel difference matrix $\Lambda_{\psi_{1},\psi_{2}}$ is sufficiently large, such that its gap with the maximum eigenvalue of the block $\Lambda_{\psi_{1},\psi_{2}}[\mathcal{I}^{c},\mathcal{I}^{c}]$ (with the complement of samples in $\mathcal{I}$ clustered by $\psi_{1}$ yet not by $\psi_{2}$) is higher than the threshold $\lambda$, then the $\mathcal{I}^{c}$-entries of the corresponding unit-norm eigenvector $\mathbf{v}$ will be bounded, reflecting the lack of $\mathcal{I}^{c}$ samples in the differentially clustered samples by embedding $\psi_{1}$ and $\psi_{2}$. Based on the above theoretical results, we propose considering the principal eigenvectors of the kernel difference matrix, and using their significant-value entries to find the subset of samples clustered by embedding $\psi_{1}$ but not grouped by $\psi_{2}$.
Since the kernel difference matrix is of size $n\times n$, a standard eigendecomposition will cost $O(n^{3})$ computations. Proposition 1 shows that the computation cost will be lower for embeddings with bounded feature maps. In fact, this result shows the computation of the eigenspace can be performed using linearly growing computation cost $O(n)$.
@@
Consider the kernel difference matrix $\Lambda_{\psi_{1},\psi_{2}}$ in (3). This matrix shares the same non-zero eigenvalues with the following matrix:
$\Gamma_{\psi_{1},\psi_{2}}=\begin{bmatrix}C_{\hskip-1.42262pt\psi_{1}}&C_{\hskip-1.42262pt\psi_{1},\psi_{2}}\vskip 5.69054pt\\ -C^{\top}_{\hskip-1.42262pt\psi_{1},\psi_{2}}&-C_{\hskip-1.42262pt\psi_{2}}\end{bmatrix}\in\mathbb{R}^{(d_{1}+d_{2})\times(d_{1}+d_{2})}$ (4)
@@
We also note that for every (right) eigenvector $\mathbf{v}\in\mathbb{R}^{d_{1}+d_{2}}$ of matrix $\Gamma_{\psi_{1},\psi_{2}}$ in (4), which we call the differential covariance matrix, we can find the corresponding eigenvector $\mathbf{u}$ of kernel difference matrix $\Lambda_{\psi_{1},\psi_{2}}$ as follows:
@@
Note that the computation of the matrix $\Gamma_{\psi_{1},\psi_{2}}$ can be performed with $O\bigl{(}n(d_{1}+d_{2})^{2}\bigr{)}$ computations, linearly growing in sample size, and the eigendecomposition of $\Gamma_{\psi_{1},\psi_{2}}$ can be handled via $O\bigl{(}(d_{1}+d_{2})^{3}\bigr{)}$, depending on the dimensions of the kernel feature maps $\phi_{1},\phi_{2}$, and finally the eigenvector mapping from $\Gamma_{\psi_{1},\psi_{2}}$ to $\Lambda_{\psi_{1},\psi_{2}}$ will be $O\bigl{(}n(d_{1}+d_{2})^{2}\bigr{)}$. Therefore, the entire eigenvector computation of $\Lambda_{\psi_{1},\psi_{2}}$ can be handled using $O\bigl{(}n(d_{1}+d_{2})^{2}+(d_{1}+d_{2})^{3}\bigr{)}$ computations. The algorithm 1 contains the main steps in computing SPEC-eigenvectors using the above approach. As detailed in this algorithm, the computation of the differential kernel covariance matrix can be run over samples in a cascade, avoiding the need to store a large dataset.
Remark 1.
Unlike the $n\times n$ kernel difference matrix $\Lambda_{\psi_{1},\psi_{2}}\in\mathbb{R}^{n\times n}$, the differential covariance matrix $\Gamma_{\psi_{1},\psi_{2}}\in\mathbb{R}^{(d_{1}+d_{2})\times(d_{1}+d_{2})}$ is not a symmetric matrix. Still, $\Gamma_{\psi_{1},\psi_{2}}$ will only have real eigenvalues, where the non-zero eigenvalues are shared with $\Lambda_{\psi_{1},\psi_{2}}$. In Appendix A.3, we further provide a Cholesky decomposition-based method to reduce the computation of $\Gamma_{\psi_{1},\psi_{2}}$’s non-zero eigenvalues and eigenvectors to the eigendecomposition of a $(d_{1}+d_{2})\times(d_{1}+d_{2})$-dimensional symmetric matrix.
Applying the standard linear and cosine-similarity kernels, the kernel feature dimension will match that of the embedding, which is usually bounded by 1000 for standard image and text embeddings. In the case of shift-invariant kernels, e.g. the Gaussian (RBF) kernel, whose feature dimension is infinite, we utilize the random Fourier features (RFFs) (Rahimi & Recht, 2007; Sutherland & Schneider, 2015) to reduce the dimension of the kernel feature dimension with a proxy kernel function characterized by the RFFs. Following the RFF framework, given a kernel function $k(x,y)=\kappa(x-y)$ that is normalized i.e. $\kappa(0)=1$, we draw a number $m$ independent Fourier features $\omega_{1},\ldots,\omega_{m}\sim\widehat{\kappa}$ from probability density function $\widehat{\kappa}$ which denotes the Fourier transform of $\kappa$ defined as
@@
Then, the RFF method approximates the shift-invariant kernel $k(x,y)\approx\widehat{k}(x,y)=\bigl{\langle}\widehat{\phi}(x),\widehat{\phi}(y)\bigr{\rangle}$ where
@@
6:  Update $C_{\psi_{1},\psi_{2}}\hskip-3.50006pt\leftarrow\hskip-3.00003ptC_{\psi_{1},\psi_{2}}\hskip-3.00003pt+\hskip-1.00006pt\frac{1}{n}\phi_{1}(\psi_{1}(\mathbf{x}_{i}))\phi_{2}(\psi_{2}(\mathbf{x}_{i}))^{\top}$